# Bringing memory to Boolean networks: a unifying framework

### v1

@@ 5 Discussion
We proposed a novel and general characterization of BN dynamics that can exploit a memory along trajectories in order to reach configurations that are not reachable with conventional (general) asynchronous updates. This led us to provide a generic framework for defining and comparing such update modes. We have shown how update modes of the literature such as the most permissive, the interval, and the cuttable modes can be expressed in this framework, emphasizing the notion of memory they employ. Besides existing modes, we took advantage of this framework to introduce novel update modes: the history-based mode, which can use any past configuration as a reference to update its current one, and the trapping mode, which can also revive and update a past configuration.
The comparison of trajectories and reachable configurations between these memory-based update modes resulted in a hierarchy of (weak) simulation, in which the trapping mode, and its equivalent subcube-based mode, subsume all others, including the most permissive mode. Indeed, while both trapping and most permissive enable to consider any past state of each component to compute the possible updates, the trapping mode allows to update a configuration from the past, making it the current one. Informally and intuitively, this could be seen as some sort of time traveling: after the update of some components, on can return to a past configuration and enable state changes based on knowledge of the future (which can then be different). Both trapping and most permissive can be characterized by the means of principal and minimal trapspaces of the BN, motivating their combinatorial study. We have notably demonstrated that trapping dynamics correspond to general asynchronous dynamics that are transitive (Prop. 4.6), and that trapping and most permissive modes coincide on the reachability of configurations in minimal trapspaces (Lemma 3.12), which in these cases, are the attractors of the dynamics. Nevertheless, we have also shown that there exists exponentially more most permissive dynamics than trapping dynamics (Prop. 3.13), the difference occurring on the reachability of transient (non-limit) configurations. Yet, most permissive dynamics being also transitive, the number of most permissive dynamics is dwarfed by the number of (a)synchronous ones.
All the studied modes have the ability to generate trajectories leading to configurations that are not reachable by applying only synchronous or asynchronous updates. These types of dynamics have been initially motivated in interval [5] and most permissive [14] as means to account for delays in the update of components. In the case of the most permissive updates, and later of the cuttable updates, these additional transitions were also motivated by the abstraction of quantitative systems, by capturing the effect of having multiple interactions depending on the quantitative state of the component: for instance, when component $i$ increases from 00 to $1$, there could be a moment before reaching $1$ whenever it is high enough to interact with component $j$ but not high enough to interact with component $j^{\prime}$.
In [15, 16], authors also provided a unifying characterization of these unconventional update modes as non-deterministic updates of components, in contrast with standard block-sequential or (a)synchronous updates, which are deterministic. These non-deterministic updates reflect some sort of strong asynchronicity within the computation of the update of one component, enabling to generate more trajectories.
@@
Here, we brought a different view on these modes, by focusing on a notion of memory used to compute next configurations from the knowledge of past ones. In addition to a complementary understanding of modes having larger dynamics than the general asynchronous, our framework can ease and motivate the definition of further update modes to be explored.

### v2

@@ 5 Discussion
We proposed a novel and unifying characterization of BN dynamics that can exploit a memory along trajectories in order to reach configurations that are not reachable with conventional (general) asynchronous updates. We have shown how update modes of the literature such as the most permissive, the interval, and the cuttable modes can be expressed in this framework, emphasizing the notion of memory they employ. Besides existing modes, we took advantage of this framework to introduce three novel update modes: the history-based mode, which can use any past configuration as a reference to update its current one, and the trapping and subcube-based modes, which can also revive and update a past configuration. The comparison of trajectories and reachable configurations between these memory-based update modes resulted in a hierarchy of (weak) simulation, in which the trapping mode, and its equivalent subcube-based mode, subsume all others, including the most permissive mode. Moreover, trapping and most permissive modes coincide on the reachability of configurations in minimal trapspaces, which in these cases, are the attractors of the dynamics.
@@
In [15, 16], authors also provided a unifying characterization of these unconventional update modes as non-deterministic updates of components, in contrast with standard block-sequential or (a)synchronous updates, which are deterministic. These non-deterministic updates reflect some sort of strong asynchronicity within the computation of the update of one component, enabling to generate more trajectories. Here, we brought a different view on these modes, by focusing on a notion of memory used to compute next configurations from the knowledge of past ones. In addition to a complementary understanding of modes having larger dynamics than the general asynchronous, our framework can ease and motivate the definition of further update modes to be explored.
The update modes and hierarchy we considered here are based on a fully asynchronous definition of the trajectories, where only one component is updated at a time. In future work, these definitions could be extended to synchronous and general asynchronous modes where all or subsets of components are updated simultaneously. Finally, the results on trapping, subcube-based, and most permissive modes call for further study on the combinatorics and structure of BN trapspaces.
@@ Acknowledgments
Appendix A Proofs of equivalence of memory-based definitions
We prove the equivalence of definitions of memory-based trajectories with the original iteration-based definitions (Sect. 2.2).
A.1 Most permissive updates
We follow the notations of Sect. 2.2.4.
Let us consider any memory-based most permissive trajectory $(x^{0},\ldots,x^{l})$. Let us define $\tilde{x}^{0}=x^{0}$, and for any $1\leq a\leq l$, $\tilde{x}^{a}\in\mathbb{P}^{n}$ such that $\Delta(\tilde{x}^{a-1},\tilde{x}^{a})=\Delta(x^{a-1},x^{a})=\{i^{a}\}$, and $\tilde{x}^{a}_{i^{a}}=\nearrow$ if $x^{a}=1$ and $\tilde{x}^{a}_{i^{a}}=\searrow$ if $x^{a}=0$. Then, remark that for any $0\leq a\leq l$, $x^{0},\ldots,x^{a}]=\gamma(\tilde{x}^{a})$. Thus, $\tilde{x}^{0}\to\ldots\to\tilde{x}^{l}$ is a valid sequence of most permissive iterations.
Conversely, let $x^{0}\to\tilde{x}^{1}\to\ldots\to\tilde{x}^{l}$ be any sequence of most permissive iterations, with $\tilde{x}^{1},\ldots,\tilde{x}^{l}\in\mathbb{P}^{n}$. Given $\tilde{x}\in\mathbb{P}^{n}$, let us define the configuration $\beta(\tilde{x})\in\mathbb{B}^{n}$ as $\beta(\tilde{x})_{i}=1$ whenever $\tilde{x}_{i}\in\{1,\nearrow\}$ and $\beta(\tilde{x})_{i}=0$ otherwise. Remark that for any $k\in\{0,\ldots,l\}$, $\gamma(\tilde{x}^{k})\subseteq[x^{0},\beta(\tilde{x}^{1}),\ldots,\beta(\tilde{x}^{k})]$. Thus, there exists a corresponding memory-based trajectory $(x^{0},\ldots,\beta(\tilde{x}^{l}))$, where we filter out iterations that change a component state from $\nearrow$ (resp. $\searrow$) to $1$ (resp. 0) to avoid successive repetition of a configuration. Formally, the corresponding memory-based trajectory is $(x^{0},\beta(\tilde{x}^{\pi^{-1}(1)}),\cdots,\beta(\tilde{x}^{\pi^{-1}(l^{\prime})}))$, with $\pi:\{0,\cdots,l\}\to\{0,\cdots,l\}$ such that $\pi(0)=0$, and for any $b\in\{1,\ldots,l\}$, $\pi(b)=\pi(b-1)$ whenever $\beta(\tilde{x}^{b})=\beta(\tilde{x}^{b-1})$ and $\pi(b)=\pi(b-1)+1$ otherwise, and $l^{\prime}=\pi(l)$.
A.2 Interval updates
We follow the notations of Sect. 2.2.6.
Let us consider any memory-based interval trajectory $(x^{0},\ldots,x^{l})$ with associated $V^{a}$ vectors and $s^{a}$ configurations for $0\leq a\leq l$, considering $s^{0}=x^{0}$. Intuitively, to obtain the corresponding interval iterations, we will map the $x^{a}$ configurations to the write nodes, and $s^{a}$ configurations to the read nodes: given any pair of configurations $x,s\in\mathbb{B}^{n}$, let us define the configuration $z(x,s)\in\mathbb{B}^{2n}$ as follows: for all $i\in[n]$, $z(x,s)_{2i-1}=x_{i}$ and $z(x,s)_{2i}=s_{i}$. Given any $1\leq a<l$, we show there is a sequence of asynchronous iterations of $\tilde{f}$ from the configuration $z(x^{a-1},s^{a-1})$ to the configuration $z(x^{a},s^{a})$. The idea is to first update the read nodes: for any component $i\in\Delta(s^{a-1},s^{a})$, by definition of $s^{a}$ and $V^{a}$, $s^{a}_{i}=x^{V^{a}_{i}}_{i}=x^{a-1}$. Therefore, $f_{2i}(z(x^{a-1},s^{a-1}))=x^{a-1}$. By updating only components in $\Delta(s^{a-1},s^{a})$, there is a sequence of iterations from $z(x^{a-1},s^{a-1})$ to $z(x^{a-1},s^{a})$. Let $i^{a}$ be the component updated between $x^{a-1}$ and $x^{a}$. Because $V^{a}_{i^{a}}=a-1$, $x^{a-1}=s^{a}$. Thus, $\tilde{f}_{2i^{a}-1}(z(x^{a-1},s^{a}))=f_{i^{a}}(s^{a})=x^{a}_{i^{a}}$. Hence, $z(x^{a-1},s^{a})\to z(x^{a},s^{a})$.
Conversely, given a sequence of interval iterations $z(x^{0},x^{0})\to z^{1}\to\ldots\to z^{l}$, one can compute a corresponding memory-based trajectory $(x^{0},x^{1}=\tau_{\text{w}}(z^{\pi^{-1}(1)}),\ldots,x^{l^{\prime}}=\tau_{\text{w}}(z^{\pi^{-1}(l^{\prime})}))$ corresponding on the projection over the write nodes, with $\tau_{\text{w}}:\mathbb{B}^{2n}\to\mathbb{B}^{n}$ maps configurations of the interval semantics to configurations of the BN $f$ by projecting on the write nodes, i.e., $\tau_{\text{w}}(z)_{i}=z_{2i-1}$ for every $i\in[n]$; and with $\pi:\{0,\cdots,l\}\to\{0,\cdots,l\}$ such that $\pi(0)=0$ and for any $b\in\{1,\ldots,l\}$, $\pi(b)=\pi(b-1)$ whenever $\tau_{\text{w}}(z^{b})=\tau_{\text{w}}(z^{b-1})$ and $\pi(b)=\pi(b-1)+1$ otherwise, and $l^{\prime}=\pi(l)$. The corresponding vectors $V^{a}$ for $1\leq a\leq l^{\prime}$ are defined such that $s^{a}_{j}=x^{V^{a}_{j}}_{j}=\tau({z^{\pi^{-1}(a)}})_{j}$ for all $j\in[n]$.
A.3 Cuttable updates
We follow the notations of Sect. 2.2.7.
The correspondence between memory-based trajectories and sequences of cuttable iterations follows the same scheme as in the previous section: between two configurations $x^{a-1}$, $x^{a}$ of a memory-based trajectory, there exists a corresponding sequence of iterations that first update the read nodes according to $C^{a}$, and then update the write node corresponding to the component $i^{a}$. Conversely, given a sequence of cuttable iterations, one obtain a corresponding memory-based trajectory by projecting over the write nodes.
Let us consider any memory-based cuttable trajectory $(x^{0},\ldots,x^{l})$ with associated $C^{a}$ matrices and $s^{a}$ configurations for $0\leq a\leq l$, considering $C^{0}=0^{n\times n}$ and $s^{0}=x^{0}$. Given a configuration $x\in\mathbb{B}^{n}$ of the BN $f$, a trajectory $(x^{0},\ldots,x^{a})$, and any matrix $C\leq(a-1)^{n\times x}$, let us define the configuration $z(x,(x^{0},\ldots,x^{a}),C)\in\mathbb{B}^{n+|E|}$ of the $E$-extension of $f$ as follows: for all $i\in[n]$, $z(x,C)_{i}=x_{i}$, and for all $(j,k)\in E$, $z(x,C)_{(j,k)}=x^{C_{k,j}}_{j}$. Given any $1\leq a<l$, we show there is a sequence of asynchronous iterations of $\tilde{f}$, the $E$-extension of $f$, from the configuration $z(x^{a-1},C^{a-1})$ to the configuration $z(x^{a},C^{a})$. The idea is to first update the read nodes, i.e., all the nodes $(j,k)\in [n]\times[n]$ such that $C^{a}_{k,j}\neq C^{a-1}_{k,j}$: by definition, $\tilde{f}_{(j,k)}(z(x^{a-1},C^{a-1}))=x_{j}^{C^{a-1}_{k,j}}$. This leads to a sequence of iterations from $z(x^{a-1},C^{a-1})$ to $z(x^{a-1},C^{a})$. Let $i^{a}$ be the component updated between $x^{a-1}$ and $x^{a}$: we obtain that $s^{a}=\pi^{i^{a}}(x^{a-1})$, thus $\tilde{f}_{i^{a}}(z(x^{a-1},C^{a}))=x^{a}_{i^{a}}$. Hence, $z(x^{a-1},C^{a})\to z(x^{a},C^{a})$.
Conversely, given a sequence of iterations $z(x^{0},C^{0})\to z^{1}\to\ldots\to z^{l}$ of the $E$-extension BN $\tilde{f}$, one can compute a corresponding memory-based trajectory $(x^{0},x^{1}=\tau_{\text{w}}(z^{\pi^{-1}(1)}),\ldots,x^{l^{\prime}}=\tau_{\text{w}}(z^{\pi^{-1}(l^{\prime})}))$ corresponding on the projection over the write nodes, with $\tau_{\text{w}}:\mathbb{B}^{n+|E|}\to\mathbb{B}^{n}$ maps configurations of $\tilde{f}$ to configurations of $f$ by projecting on the write nodes, i.e., $\tau_{\text{w}}(z)_{i}=z_{i}$ for every $i\in[n]$; and with $\pi:\{0,\cdots,l\}\to\{0,\cdots,l\}$ such that $\pi(0)=0$ and for any $b\in\{1,\ldots,l\}$, $\pi(b)=\pi(b-1)$ whenever $\tau_{\text{w}}(z^{b})=\tau_{\text{w}}(z^{b-1})$ and $\pi(b)=\pi(b-1)+1$ otherwise, and $l^{\prime}=\pi(l)$. The corresponding matrices $C^{a}$ for $1\leq a\leq l^{\prime}$ are defined such that $x^{C^{a}_{k,j}}_{j}=z^{\pi^{-1}(a)}_{(j,k)}$ for all $j,k\in[n]$.